OpenStudy (anonymous):

Solving Rational Functions 2/x+3/5=22/5x HELP!!!

OpenStudy (owlcoffee):

Does it look like this?: \[\frac{ \frac{ 2 }{ x+3 } }{ 5 }=\frac{ 22 }{ 5x }\]